# Crosslingual Capabilities and Knowledge Barriers in Multilingual Large Language Models

### v1

@@ 3.1 Crosslingual knowledge barrier in general knowledge
Figure 4: Examples of original, full-translated, and proposed mixed-language multiple choice question (MCQ) formats on MMLU.
@@
In the above setups, even if a model has independently acquired knowledge in multiple languages, it must rely on crosslingual capabilities to select the correct answer. These tasks pose additional difficulty because, generally, the answer is expected to be in the same language as the questions. We use Google Translate API for all translations, and all derived datasets have the same size as the original dataset.
Open-source models are evaluated in the standard way by calculating the likelihood of each option token and using the maximum one as the model prediction (Touvron et al., 2023). For the close-soured models, we directly use the predicted best option (i.e., first token) by the model as answer.
@@
Crosslingual barrier in MMLU knowledge. We focus on the MMLU benchmark for evaluating general knowledge, which comprises 4-option MCQs and includes $14k$ test samples from 4 domain categories (i.e., STEM, Social Science, Humanities, Others) across 57 subjects. The diversity of these domains enables us to draw general observations. (1) The monolingual evaluation results in Figure 3 show that all LLMs consistently achieve higher accuracy when MCQs are presented in English compared to other languages. This is likely because the relevant general knowledge is more frequently presented in English within the pretraining corpus, and LLMs struggle to transfer this knowledge to other languages automatically. Additionally, performance in German (de) is worse across all tasks and models, aligning with existing evaluations (Mistral, 2024), where Mistral and Llama models show lower accuracy in German than in other languages on commonsense reasoning benchmarks. Moreover, LLMs perform poorly in translation from English to German (Table 1), suggesting that the models may have inferior multilingual capabilities in German. (2) The results in Figure 5 demonstrate a notable accuracy drop in the mixed-language settings, including question+GT-option, GT-option, and mixup translations, compared to monolingual settings (i.e., English and full-translation). This suggests that LLMs struggle to understand the more difficult contexts in multiple languages and to relate the corresponding knowledge effectively to answer MCQs, highlighting a crosslingual knowledge barrier in the MMLU benchmark. We note such barrier exists even for the state-of-the-art models like GPT-4 (e.g., $68.6181.82\xrightarrow{}$ 68.61 when comparing English to mixup-translated MMLU). (3) The GT-option translation setting leads to the worst performance, indicating an inherent behavioral bias of LLMs that tends to avoid selecting a non-English option, even if it is the correct choice. This bias is further supported by the controlled comparisons in one-wrong-option translation settings, where LLMs achieve even higher accuracy than the English setting, as the model leverages the bias and avoids selecting the (incorrect) non-English option. (4) LLMs obtain higher accuracy on question-translated and options-translated settings than full-translated settings, likely because the former two settings still have remaining context in English, which helps the models perform better.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Accuracy on MMLU variant benchmarks. The bars with * denotes the average accuracy across fr, de, es, and it. LLMs perform better at answering MCQs in English than in mixed-language settings, especially the ground truth option and mixup translation, indicating the existence of cross-lingual knowledge barriers. Due to budget constraints, GPT-3.5 and GPT-4 are evaluated only in the most challenging settings.
3.2 Crosslingual knowledge barrier in domain-specific knowledge
@@
Harry Potter Quiz. We use the Harry Potter world for the domain-specific knowledge evaluations, as it revolves around a highly detailed and extensive fictional universe with its own unique characters, terminology, and concepts. We manually curate a multiple-choice question-answering dataset called the Harry Potter Quiz (HP-Quiz) by extracting information from the Harry Potter Wiki pages [^4] https://harrypotter.fandom.com/wiki/Main_Page. Further details about the dataset are provided in Appendix A.
@@
Evaluation. Each multiple-choice question in the HP-Quiz dataset is available in five different languages: English (en), French (fr), German (de), Spanish (es), and Italian (it). To assess the crosslingual knowledge barrier, we consider both (1) the original model, and (2) the model fine-tuned on domain-specific corpora [^5] Specifically, we preprocess the WikiText-103 dataset (Merity et al., 2017) and select documents highly relevant to the Harry Potter universe using a retriever (see Section B.3 for details). presented only in English. For evaluation, we prompt the model with the multiple-choice question in each language, and report the accuracy of the model in selecting the correct answer in this multiple-choice task for each language.
Figure 6: Multiple-choice accuracy of various multilingual LLMs on the Harry Potter Quiz benchmark before (left) and after (right) fine-tuning the model on in-domain content presented in English (i.e., Harry Potter related documents selected from WikiText-103). Models consistently perform better at answering questions in English than other languages, indicating the presence of a crosslingual knowledge barrier. Notably, for Mistral, fine-tuning results in decreased performance on the HP-Quiz, potentially due to a mismatch between the fine-tuning data mixture and the pretraining data mixture, a phenomenon also observed by Xia et al. (2024).
@@
Crosslingual barrier also exists for Harry Potter knowledge. As shown in Figure 6, when presented with the same set of questions in 5 languages, the model consistently exhibits higher accuracy in answering correctly in English. This trend holds for both pretrained LLMs (left) and fine-tuned LLMs (right). After fine-tuning on domain-specific English corpora, despite the increase in model accuracy in English (which is more evident for Llama2-7B and Llama3-8B), the crosslingual knowledge barrier persists. These observations provide compelling evidence that the crosslingual knowledge barrier extends beyond general knowledge into specific domains.
4 Overcoming crosslingual knowledge barriers
@@
In this section, we explore potential methods to overcome the manifest crosslingual knowledge barrier that we identified in the existing multilingual LLMs. We consider two types of potential mitigation methods, including inference-time intervention (Section 4.1), including prompt engineering and few-shot demonstration, and training-time intervention (Section 4.2), including mixed-language fine-tuning on general and domain-specific corpora.
4.1 Inference-time mitigation falls short of overcoming crosslingual knowledge barriers
@@
We evaluate inference-time mitigation methods to improve LLM performance on the mixup-translated MMLU, a challenging crosslingual setting evidenced by the low performance in Figure 5.
@@
Prompt engineering. We evaluate the following prompting strategies: (1) Alternative option ID characters. We replace the default A/B/C/D with a/b/c/d or 1/2/3/4, motivated by recent evidence on selection bias in option IDs for MCQ tasks (Zheng et al., 2024) and to account for the possibility that the Arabic numerals are more invariant to languages. (2) Multilingual awareness instruction: We add an explicit instruction before the MCQs (e.g., “Remember that the question and options can be in different languages”) to make models aware of the potential presence of other languages.
@@
Few-shot demonstration. Our evaluation mainly considers the 0-shot setting, which excludes any biases introduced by the few-shot demonstrations (Zhao et al., 2021), but we also conduct 5-shot experiments to further investigate crosslingual performance. MMLU covers 57 subjects, and the few-shot demonstrations for each subject are derived from the corresponding development set and shared across all test samples within the same subject. We employ several strategies to construct few-shot demonstrations: (1) English demonstration: English MCQ and answer pairs. (2) Same bias demonstration: mixup-translated MCQ and answer pairs, where each MCQ demonstration is constructed in the same way as the test sample. (3) Translate-Then-Answer demonstration: For each mixup-translated MCQ, we prompt LLMs first to translate it into English before producing the final answer. To help LLMs follow the explicit translation instruction, we provide demonstrations where each includes a mixup-translated MCQ, the corresponding English MCQ, and its answer. We provide the detailed prompt templates in Section B.2.
@@
Mixup (5-shot) English demonstrations $35.23$ $43.15$ $49.46$ $50.99$
Same bias demonstrations $36.92$downarrow4.614.614.614.61 $44.32$downarrow7.797.797.797.79 $51.07$downarrow9.149.149.149.14 $51.65$downarrow8.898.898.898.89
Translate-Then-Answer demonstrations $30.02$ $42.93$ $42.27$ $47.79$
From the results in Table 3, (1) regarding prompt engineering, we observe no improvement and even a performance drop compared to the default prompt. It suggests that the crosslingual knowledge barrier is an inherent failure of LLMs that cannot be effectively addressed by simple prompt engineering. (2) 5-shot settings consistently improve performance compared to 0-shot settings on mixup MMLU because providing demonstrations in the corresponding subject helps LLMs generalize to knowledge-intensive tasks. (3) Mixup demonstrations lead to better performance than English demonstrations because the mixed language pattern in the demonstrations matches that of the test examples. (4) Translate-Then-Answer demonstrations are not effective. We observe failure patterns where, after translating to English, sometimes LLMs merely continue generating text without outputting the desired answer for the MCQ task. (5) Even under the best demonstration strategy, there still exhibits a substantial accuracy gap in mixup MMLU compared to English MMLU. Consequently, we explore training-time intervention in Section 4.2 via mixed language fine-tuning methods.
4.2 Mixed-language fine-tuning effectively mitigates crosslingual knowledge barriers
Given the limited success of inference-time interventions, we turn our attention to training-based methods that aim to directly instill better crosslingual knowledge in the model itself. Specifically, we explore mixed-language fine-tuning, where we explicitly construct a fine-tuning dataset comprising examples from multiple languages. To ensure a balanced representation of different languages, we split the training data into smaller units and randomly select a target language for each unit, translating the unit into that language if necessary. This approach also ensures that the translated data is of similar size as the original English data, enabling a fair comparison. Note that this approach differs from using parallel corpora as each unit is only presented in a single language.
We explore different choices for the smallest unit for translation, including the following settings:
  • Full document translation: the entire document (example) is translated to a random language.

### v2

@@ 3 Identify crosslingual knowledge barrier
(a) Examples of original, full-translated, and proposed mixed-language Multi-choice Question (MCQ) formats.
(b) Monolingual eval.
(c) Crosslingual eval.
(d) Crosslingual results for each lang.
Figure 3: (a) presents examples of original, full-translated, and proposed mixed-language Multi-choice Question (MCQ) formats. (b) shows the monolingual evaluation under 5 languages where 15 LLMs all perform better at answering MMLU MCQs in English. Detailed results under four MMLU domains (STEM, Social Science, Humanities, Others) are in Fig. 12. (c) shows the results under cross-lingual settings, where * denote the average accuracy across {fr, de, es, it}. LLMs perform worse at answering MCQs in mixed-language settings than in English, especially the GT-option and Mixup translation, indicating the existence of cross-lingual knowledge barriers. (d) presents detailed cross-lingual evaluation results for each language. We observe similar findings for all 15 LLMs in Fig. 13.
While multilingual LLMs have demonstrated impressive explicit crosslingual abilities, e.g., performing translations for source sequence given in the context, questions remain about their capability to implicitly retrieve and utilize parametric knowledge across languages. For example, the model might be asked a question in one language (e.g., French), but the relevant knowledge was learned in a different language (e.g., English). As we will show in this section, LLMs struggle to seamlessly bridge the language gap when faced with tasks demanding implicit crosslingual knowledge transfer. We term this phenomenon the crosslingual knowledge barrier. In the following, we demonstrate the presence of such barriers for both general knowledge (§ 3.1) acquired during pretraining and domain-specific knowledge (§ 3.2) obtained through explicit fine-tuning.
3.1 Crosslingual barrier in general knowledge
Setup. (1) Data: We focus on the MMLU benchmark for evaluating general knowledge, which comprises 4-option Multiple Choice Question (MCQ) and includes $14k$ test samples from 4 domain categories (i.e., STEM, Social Sciences, Humanities, Others) across 57 subjects. The diversity of these domains enables us to draw general observations. As in standard MCQ evaluation (Touvron et al., 2023; Zheng et al., 2024), we do the following: for open-source LLMs, we calculate the likelihood of each option token (e.g., A/B/C/D) and using the maximum one as model prediction. For closed-source LLMs where token likelihoods are not accessible, we use the predicted best option (i.e., first token) with decoding temperature 00 as answer. We additionally compare these two evaluation strategies on open-source LLMs in Tb. 8. (2) Models: In addition to the six popular LLMs mentioned in § 2, we evaluate nine additional LLMs, including strong multilingual model Aya-23-8B, Aya-expanse-8b (Aryabumi et al., 2024), Llama-3.1-8B, Qwen2.5-7B, Mistral-Nemo, two Tower-series models trained under cross-lingual supervision, and two models beyond traditional Transformers—Zamba-7B, a state-space model (Glorioso et al., 2024), and Mistral-8x7B, a Mixture-of-Experts model (Jiang et al., 2024).
Monolingual evaluation is inadequate for assessing crosslingual abilities. Previous studies have evaluated MCQ tasks on general knowledge in multilingual settings. For instance, Mistral series models (Mistral, 2024) were benchmarked on translated versions of MMLU (Hendrycks et al., 2021) dataset in French (fr), German (de), Spanish (es), and Italian (it), separately. We refer to such monolingual evaluation setup as “full-translation”. We report the monolingual evaluation results in Fig. 3(b). While such results indicate multilingual proficiency, they are insufficient to show crosslingual proficiency. This is because the relevant general knowledge might be present in each of the evaluated languages in the pretraining dataset, so the LLMs could answer the full-translated questions based on knowledge learned in each individual language without invoking crosslingual capabilities. Such possibility is difficult to verify, as pretraining data for most LLMs are undisclosed.
Mixed-language evaluation. To directly invoke crosslingual capabilities of LLMs we suggest a mixed-language MCQ formats. The motivation is that the models are less likely exposed to such format during training, while a truly crosslingual agent like a human speaking multiple languages would be able to answer them as well as the monolingual versions. Specifically, we propose the following formats purposefully designed to be novel compositions unlikely to have been encountered in pretraining (examples in Fig. 3(a)): (1) Mixup trans: translating question and all options into 5 different languages, with the language assignments randomly determined from {en, fr, de, es, it}. (2) Question trans: translating question into a non-English language. (3) Options trans: translating all options into a non-English language. (4) Question+GT-option trans: translating both question and ground truth option into a non-English language, while keeping the remaining options in English. (5) GT-option trans: translating ground truth option into a non-English language, while keeping question and the rest of the options in English. (6) One-wrong-option trans: randomly selecting one incorrect option and translating it into a non-English language.
@@ 3.1 Crosslingual barrier in general knowledge
In above settings, even if a model has independently acquired knowledge in multiple languages, it will have to rely on crosslingual capabilities to select the correct answer. We perform translation via the Google Translate API, and all derived datasets have the same size as the original one.
Figure 4: Crosslingual knowledge barriers across 16 languages under mixed-language MCQ evaluation in MMLU knowledge.
@@
Crosslingual barrier in MMLU knowledge in 15 LLMs. (1) The results in Fig. 3(c) and Fig. 3(d) demonstrate a notable accuracy drop in the mixed-language settings, including Question+GT-option, GT-option, and Mixup translations, compared to monolingual settings (i.e., English and full-translation). This suggests that LLMs struggle to understand the more difficult contexts in multiple languages and to relate the corresponding parametric knowledge effectively to answer MCQs, highlighting a crosslingual knowledge barrier in the MMLU benchmark. We note such barrier exists even for the state-of-the-art models like GPT-4 (e.g., $68.6181.82\xrightarrow{}$ 68.61 when comparing English to mixup-translated MMLU). (2) The GT-option translation setting leads to the worst performance, indicating an inherent behavioral bias of LLMs that tends to avoid selecting a non-English option, even if it is the correct choice. This bias is further supported by the controlled comparisons in One-wrong-option translation settings, where LLMs achieve even higher accuracy than the English setting, as the model leverages the bias and avoids selecting the (incorrect) non-English option. (3) LLMs obtain higher accuracy on Question-translated and Options-translated settings than Full-translated settings, likely because the MCQs under the former two settings still have remaining context in English, which helps the models perform better.
Evaluation on 16 languages. To demonstrate the universality of our findings, we evaluate 11 additonal languages: (1) Low-resource languages (Zhang et al., 2023c): Malay (ms), Danish (da), Finnish (fi), Norwegian (no), Bengali (bn), Amharic (am); (2) Languages with token distributions significantly different from English: Russian (ru), Chinese (zn), Hebrew (he), Arbic (ar) and Hindi (hi). Fig. 4 shows the performance gaps between English (dashed line) and other languages persist in both monolingual (Full translation) and mixed-language (Options/GT-option translation) settings. This gap is particularly evident for low-resource languages like Finnish (fi), Bengali (bn), and Amharic (am), revealing the general challenge of cross-lingual knowledge barriers. Moreover, Llama-3.1-8B has a more balanced performance across languages than Qwen2.5-7B and Aya-expanse-8B. For most non-English languages, multilingual LLM show the weakest performance when ground-truth options require cross-lingual reasoning.
3.2 Crosslingual barrier in domain knowledge
In § 3.1, we demonstrated the crosslingual knowledge barrier for off-the-shelf LLMs in general knowledge required to solve MMLU tasks, where we assume this knowledge was obtained during pretraining. Here, we present a more controlled test through explicit fine-tuning on domain-specific knowledge, aiming to answer the following question: Could the model utilize the domain-specific knowledge (e.g., Harry Potter facts) acquired in one language (e.g., English) via fine-tuning to answer questions about this knowledge in other languages? We will show that crosslingual knowledge barrier also exists for domain-specific knowledge.
@@ 3.2 Crosslingual barrier in domain knowledge
Domain-specific datasets: (1) Harry Potter Quiz. We use the Harry Potter world for evaluations, as it revolves around a highly detailed and extensive fictional universe with its own unique characters, terminology, and concepts. We manually curate a multiple-choice question-answering dataset called the Harry Potter Quiz (HP-Quiz) by extracting information from the Harry Potter Wiki pages [^3] https://harrypotter.fandom.com/wiki/Main_Page. Further details about the dataset are provided in § C. (2) TOFU (Maini et al., 2024) introduces synthetic knowledge absent from pretrained LLMs’ training data. Specifically, the TOFU dataset contains question-answer pairs derived from fictional autobiographies of 200 non-existent authors, generated by GPT-4 using predefined author attributes.
@@
English Fine-tuning. To assess the cross-lingual knowledge barrier, we consider both (a) original LLMs, and (b) LLMs fine-tuned on domain-specific corpora presented only in English. Specifically, we preprocess the WikiText-103 dataset (Merity et al., 2017) and select documents highly relevant to the Harry Potter universe using a retriever (see § D.3 for details) as our fine-tuning data, and evaluate on HP-quiz. For TOFU, we use the original dataset for English fine-tuning, and the paraphrased questions for evaluation.
Evaluation. For HP-Quiz, we evaluate the model by prompting it with multiple-choice questions in each language and report its accuracy in selecting the correct option (i.e., A/B/C/D). For TOFU, we prompt the model with questions in each language and compute ROUGE scores between the model’s response and the ground-truth English answer (used due to LLM’s tendency to output English after English fine-tuning, instead of the language in question).
(a) MCQ accuracy of LLMs on the Harry Potter Quiz, before (left) and after (right) fine-tuning the model on in-domain content in English (i.e., HP-related documents selected from WikiText-103).
(b) ROUGE scores of LLMs on open-ended TOFU QA tasks, before (left) and after (right) fine-tuning the model on the TOFU dataset (Maini et al., 2024) in English.
Figure 5: Models consistently perform best at answering questions in English, both before and after fine-tuning, indicating the presence of a crosslingual knowledge barrier for domain-specific Harry Potter knowledge (a) and TOFU knowledge (b).
@@
Crosslingual barrier also exists for Harry Potter and TOFU knowledge. As shown in Fig. 5(a) for HP-Quiz and Fig. 5(b) for TOFU, when presented with the same set of questions in 5 languages, the model consistently exhibits higher accuracy in answering correctly in English. This trend holds for both pretrained & fine-tuned LLMs. After fine-tuning on domain-specific English corpora, despite the increase in model accuracy in English, the crosslingual knowledge barrier persists. This suggests that LLMs struggle to fully utilize the parametric knowledge acquired during English fine-tuning to answer related questions in other languages, which shows that the crosslingual knowledge barrier extends beyond general knowledge into specific domains.
As HP represents a widely known knowledge domain with practical relevance (e.g., users may query LLMs about HP in their native languages to understand the books better), we evaluate 11 off-the-shelf LLMs on HP Quiz across 16 languages in Fig. 6. Results show LLMs perform best in English, with significantly lower accuracy on low-resource languages (e.g., bn, am) compared to high-resource ones (e.g., fr, de, es) which are our main focus. Notably, Mistral-7B and Llama3-8B achieve competitive performance against multilingual-focused models like the Aya and Tower series, validating our model and language selection in Fig. 5(a).
Figure 6: Accuracy of pretrained LLMs on HP-Quiz across 16 languages. Models perform best in English and perform better in high-resource languages (e.g., fr, de, es) than in low-resource ones (e.g., bn, am). Notably, Mistral-7B-v0.1 and Llama-3-8B show competitive performance compared to multilingual-focused models such as the Aya series and Tower series.
4 Overcome crosslingual knowledge barriers
@@
In this section, we explore potential methods to overcome crosslingual knowledge barrier we identified in multilingual LLMs. We consider two types of potential mitigation methods, inference-time interventions (§ 4.1), including prompt engineering and few-shot demonstrations, and training-time interventions (§ 4.2), including mixed-language fine-tuning on general and domain-specific corpora.
4.1 Inference-time mitigation
@@
We study inference-time mitigation to improve LLM performance on mixup-translated MMLU, a challenging crosslingual setting evidenced by the low performance in Fig. 3(c).
@@
Prompt engineering. We evaluate the following prompting strategies: (1) Alternative option ID characters. We replace the default A/B/C/D with a/b/c/d or 1/2/3/4, motivated by recent evidence on selection bias in option IDs for MCQ tasks (Zheng et al., 2024) and to account for the possibility that the Arabic numerals are more invariant to languages. (2) Multilingual awareness instruction: We add an explicit instruction before MCQs (e.g., “Remember that the question and options can be in different languages”) to make LLMs aware of the potential presence of other languages.
@@
Few-shot demonstrations. Our evaluation mainly considers the 0-shot setting, which excludes any biases introduced by the few-shot demonstrations (Zhao et al., 2021), but we also conduct 5-shot experiments to further investigate crosslingual performance. MMLU covers 57 subjects, and the few-shot demonstrations for each subject are derived from the corresponding development set and shared across all test samples within the same subject. We employ several strategies to construct few-shot demonstrations: (1) English demonstration: English MCQ and answer pairs. (2) Same bias demonstration: mixup-translated MCQ and answer pairs, where each MCQ demonstration is constructed in the same way as the test sample. (3) Translate-then-Answer demonstration: For each mixup-translated MCQ, we prompt LLMs first to translate it into English before producing the final answer. To help LLMs follow the explicit translation instruction, we provide demonstrations where each includes a mixup-translated MCQ, the corresponding English MCQ, and its answer. We provide the detailed prompts in § D.2.
@@
Mixup (5-shot) English demos $35.23$ $43.15$ $49.46$ $50.99$
Same bias demos $36.92$downarrow4.614.614.614.61 $44.32$downarrow7.797.797.797.79 $51.07$downarrow9.149.149.149.14 $51.65$downarrow8.898.898.898.89
Translate-then-Answer demos $30.02$ $42.93$ $42.27$ $47.79$
From results in Tb. 2, (1) regarding prompt engineering, we observe no improvement and even a performance drop compared to the default prompt. It suggests that the crosslingual knowledge barrier is an inherent failure of LLMs that cannot be effectively addressed by simple prompt engineering. (2) 5-shot settings consistently improve performance compared to 0-shot settings on mixup MMLU, as providing demonstrations in the corresponding subject helps LLMs generalize to knowledge-intensive tasks. (3) Mixup demonstrations lead to better performance than English demonstrations because the mixed language pattern in the demonstrations matches that of the test examples. (4) Translate-then-Answer demonstrations are not effective. We observe failure patterns where, after translating to English, sometimes LLMs merely continue generating text without outputting the desired answer for the MCQ task. (5) Even under the best demonstration strategy, there still exists a substantial accuracy gap in mixup MMLU compared to English MMLU.
4.2 Mixed-language fine-tuning
Given the limited success of inference-time interventions, we turn our attention to training-based methods that directly instill better crosslingual knowledge in the model itself.
Proposed method. Specifically, we explore mixed-language fine-tuning (FT), where we explicitly construct a fine-tuning dataset comprising examples from multiple languages. To ensure a balanced representation of different languages, we split the training data into smaller units and randomly select a target language for each unit, translating the unit into that language if necessary. This approach also ensures that the translated data is of similar size as the original English data, enabling a fair comparison. Note that this approach differs from using parallel corpora as each unit is only presented in a single language.